OpenStudy (beststudent):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. (y + 1) = 3(x – 2) B. (y – 1) = 3(x + 2) C. (x + 2) = 3(y – 1) D. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

B ><?

OpenStudy (anonymous):

correct. if the line has an equation of y = mx + c, then the slope is m if the line passes through the point (a, b), you can substitute x = a and y = b into the equation

OpenStudy (texaschic101):

y - y1 = m(x - x1)......with slope(m) = 3 and points (-2,1) y - 1 = 3(x - (-2) y - 1 = 3(x + 2) correct...it is B.